OpenStudy (anonymous):

√(ab+c^2) where a=4, b=-2 and c=-3

OpenStudy (amistre64):

thats just plug and play; the math itself is self-evident

OpenStudy (amistre64):

√(4(-2) + (-3)^2)

OpenStudy (anonymous):

\[\sqrt{(4)(-2) + (-3)^2}=\sqrt{1} = 1\]

OpenStudy (anonymous):

try using the website web2.0calc its a free scientific calculator